You have been given the following information about the production of Usher Co., and are asked to provide the plant manager with information for a meeting with the vice president of operations.

Standard Cost Card
Direct materials (8 pounds at $5 per pound) $40.00
Direct labor (0.70 hours at $10) 7.00
Variable overhead (0.70 hours at $3 per hour) 2.10
Fixed overhead (0.70 hours at $7 per hour) 4.90
$54.00


The following is a variance report for the most recent period of operations.

Variances

Costs

Total Standard Cost

Price

Quantity

Direct materials $412,000 $8,434 F $9,700 U
Direct labor 72,100 4,704 U 6,300 U

How many units were produced during the period?

Number of units

What was the actual cost per pound of raw materials? (Round answer to 2 decimal places, e.g. 1.25.)

Actual cost per pound of raw materials

$

/lb

How many actual direct labor hours were worked during the period?

Actual direct labor hours hours



What was the actual rate paid per direct labor hour? (Round answer to 2 decimal places, e.g. 1.25.)

Actual rate paid per direct labor hour

$

/hr

How many pounds of raw materials were purchased and used during the period?

Raw material pounds
Answer #1

1) Number of units produced = 412000/40 = 10300 Units

b) Material Quantity variance = (SQ-AQ)SP

-9700 = (10300*8*5-5X)

5X = 412000+9700

X(actual quantity) = 84340 Pounds

Material price variance = (SP-AP)AQ

8434 = (5*84340-X*84340)

84340X = 413266

X(Actual cost per pound) = 4.9

c) Labor quantity variance = (SH-AH)SR

-6300 = (10300*.7*10-10X)

10X = 78400

X(Actual hour) = 7840 Hour

d) Labor rate variance = (SR-AR)SH

-4704 = (10*7840-7840X)

7840X = 83104

X(Actual rate) = 10.60

e)

Material Quantity variance = (SQ-AQ)SP

-9700 = (10300*8*5-5X)

5X = 412000+9700

X(actual quantity) = 84340 Pounds